OpenStudy (kingj):

A monument is built out of toothpicks. There are 35 rows and each row has six more toothpicks than the row above. The top row has five toothpicks. Part 1: How many toothpicks are on the bottom row? Part 2: Using complete sentences, explain the procedure taken to answer this question.

OpenStudy (mertsj):

Aren't you just asking for the 35th term of the sequence whose first term is 65 and whose common difference is 6?

OpenStudy (mertsj):

\[a _{n}=a _{1}+(n-1)d\]

OpenStudy (anonymous):

a=5 d=6 n=35

OpenStudy (mertsj):

Can you plug in this time?

OpenStudy (anonymous):

\[a _{35}=5+(35-1)d\]

OpenStudy (anonymous):

5+34*6=5+204=209